Morgan Stanley has a current cash flow (at time 0) of $3.4 m and pays no dividends. The present value of the company’s future cash flows is $14.6 m. The firm is entirely financed with equity and has 400,000 shares outstanding. Assume the dividend tax rate is zero.

a.   What is the share price of Morgan Stanley stock?

b.   Suppose the board of directors of Morgan Stanley announces its plan to payout 40% of its current cash flow as dividends to its shareholders. How can Andy, who owns 800 shares of Morgan Stanley stock, achieve a zero payout policy on his own?

 



1
Expert's answer
2022-02-13T11:48:11-0500

a.DCF=3.4+14.6=18

"the share price=\frac{18 000 000}{400 000}=45"


b.may be company is having potential investment projects with positive NPV.